OpenStudy (emac):

Find the average rate of change 50x/x+3 on (2,7)

OpenStudy (3mar):

If I may help?

OpenStudy (3mar):

@emac

OpenStudy (3mar):

your equation is \[y=\frac{ 50x }{ x+3 }\] The rate of change of a function is the first derivative of this function \[\frac{ dy }{ dx }=y'=y=\frac{ (x+3)*[50]-50x(1) }{ (x+3)^2 }=\frac{ 150 }{ (x+3)^2 }\]

OpenStudy (3mar):

To find the derivative at (2,7), substitute with x=2: \[y'=\frac{ 150 }{ (2+3)^2 }=\frac{ 150 }{ (5)^2 }=\frac{ 150 }{ 25 }=6\] That is it! I hope that helps!

OpenStudy (emac):

yes thank you!!

OpenStudy (3mar):

You are welcome! Any Help... Any Time...

OpenStudy (3mar):

Thank you for the medal!